The sum of 3 numbers is 2009. If the ratio of the first number to the second is 2:3 and the ratio of the second to the third number is 5:8, what is the third number?
Answer by richwmiller(17219)

x+y+z=2009,
x/y=2/3,
y/z=5/8
x = 410, y = 615, z = 984